OpenStudy (anonymous):

Solve the equation By factoring. x^2+7x = -12 A. x = 2; x = 6 B.x = 3; x = 4 C.x = -3; x = -4 D.x = -1; x = 12

OpenStudy (anonymous):

B

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

yup, do you need to know how i got it?

OpenStudy (anonymous):

\[x^{2}+7x+12=0\] 1*12=12 4*3=12 4+3=7 write 7x=4x+3x and then make factors.

OpenStudy (anonymous):

no but thanks :)